OpenStudy (anonymous):

x/3 - 10 = -12 Solve The equation please helppppppp

OpenStudy (anonymous):

what is the x/3 ???

OpenStudy (anonymous):

it is x/3 but like x on top 3 on bottom

OpenStudy (anonymous):

ok im not sure what sign it is but ill give it a try

OpenStudy (anonymous):

i think it is x divided 3

OpenStudy (anonymous):

well i got x=1

OpenStudy (anonymous):

how

OpenStudy (anonymous):

did u get 1

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3x-10=12 +10 +10 3x=22 x=1

OpenStudy (anonymous):

but i may be wrong

OpenStudy (anonymous):

i think its x = -6

OpenStudy (anonymous):

3 goes into - 6 -2 times - 10 = -12

OpenStudy (anonymous):

were you doing x(3-10)=12 ?